OpenStudy (anonymous):

You place a cup of 210oF coffee on a table in a room that is 68oF, and 10 minutes later, it is 200oF. Approximately how long will it be before the coffee is 180oF?

OpenStudy (anonymous):

can anyone help me using newton's law of cooling

OpenStudy (dumbcow):

\[\frac{dT}{dt} = k (T - 68)\] \[\int\limits \frac{dT}{T - 68} = k \int\limits dt\] \[\ln (T-68) = kt + C\] \[T = C e^{kt} + 68\] Given: t=0, T = 210 t = 10, T = 200 \[\rightarrow C = 142\] \[\rightarrow k = \frac{1}{10}\ln (\frac{132}{142})\] Find t when T = 180